OpenStudy (katielong):

have i done this right? solve this simultaneous equation: 4x+y=1 2x-3y=18 4x-6y=36 4x+y=1 (-) -7y=35 -y=5 y=-5 thanks =)

OpenStudy (anonymous):

what about x? y is correct

OpenStudy (katielong):

1.5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x+y =1 --> y = 1 - 4x 2x-3y=18 => 2x -3 (1 -4x) = 18 2x - 3 + 12x = 18 14x = 21 ->x = 21/14 = 3/2 => y = 1 - (4 * 3/2) = 1 - 6 = -5 (3/2 , -5)

OpenStudy (katielong):

i perfer my way...